OpenStudy (anonymous):

Eighty ounces of Lazy Lawn fertilizer covers 1,250 square feet of lawn. a. How many ounces would be required to cover a 4,000-square-foot lawn? b. If Lazy Lawn costs $1.19 for a 32-ounce bag, what is the total cost to fertilize the lawn?

OpenStudy (anonymous):

It's a proportion. 80/1250=x/4000 x = (4000*80)/(1250) = 256

OpenStudy (anonymous):

Salim sells fax machines that come in Standard and Enhanced models. The Standard model sells for $325 and the Enhanced model sells for $385. If Salim sold a total of 35 units and took in $12,575, how many of the Enhanced model did he sell?

OpenStudy (anonymous):

Sx+Ey=12575 x+y=35 Here's a starting point. No full answers this time.